Consider a soot particle of mass msoot = 1.0 nanogram (=1.0 x 10-9 g = 1.0 x 10-12 kg), which corresponds to a diameter of a few microns. Some number of electrons have been added to give the particle a total charge qsoot, suppose the collector has a total charge of qcollector = 1.0 x 10-6 C and is small enough to be treated as a point charge at the rim of the smokestack. (a) if the separation between the collector and the soot particle is r = 0.10 m, what is the value of qsoot so that the electric force exerted on the particle is equal to its weight? (b) How many electrons must be added to the soot particle?



1
Expert's answer
2022-03-21T17:57:23-0400

a)

"\frac{kq_cq_s}{r^2}=mg,"

"q_s=\frac{mgr^2}{kq_c}=1.09\cdot 10^{-17}~C,"

b)

"n=\frac{q_s}e=68."